rEsTiNpEaCe:

What is ΔE for the transition of an electron from n = 8 to n = 4 in a hydrogen atom? What is the frequency, ν, of the spectral line produced? Enter your answers with 4 significant digits. Express your answer for ΔE in J Express your answer for ν in s−1.

rEsTiNpEaCe:

@dude

rEsTiNpEaCe:

@Tranquility

rEsTiNpEaCe:

I got: ∆E = 1.020 x 10^-19 J v = 1.539 x 10^14 s^-1 But it was wrong. Can you help?

rEsTiNpEaCe:

∆E = (-13.6/8^2 - (-13.6/4^2) ) x 1.6x10^-19 = 1.02 x 10^-19 J I used the Bohr model for energy to find delta ∆E and used 1.6x10^-19 to convert to Joules. v = ∆E/h = 1.02 x 10^-19/6.626x10^-34 = 1.5393 x 10^14 s^-1

rEsTiNpEaCe:

@Vocaloid

rEsTiNpEaCe:

My v value is correct, but it says my ∆E value isn't. I think it has to do with something with the four significant digits

Tranquility:

((-13.6/(8^2) - (-13.6/4^2))) * 1.6022*10^-19 = \(\boxed{1.021\times 10^{-19}}\) 1 eV = 1.6022*10-19 J maybe using 1.6022 instead of 1.6

rEsTiNpEaCe:

I ended up figuring it out, it was that it was -1.021 since energy was lost. but thanks!

Tranquility:

Glad you figured it out! It's interesting that the - sign is required since we're calculating the change (\(\Delta\)) in energy